OpenStudy (anonymous):

What is cos 0 if sec 0 = 2?

Nnesha (nnesha):

what's reciprocal of cos ?

OpenStudy (anonymous):

1?

Nnesha (nnesha):

nope like tan = sin /cos sin = 1/csc cos = ?

OpenStudy (anonymous):

1/cos idk :(

Nnesha (nnesha):

cos = 1/sec sec = 1/cos

Nnesha (nnesha):

cos = 1/ sec sec =2 replace sec by 2 that's it

OpenStudy (anonymous):

thank you :)

Nnesha (nnesha):

np :-)